You split $2000 between two savings accounts. Account A pays 2% annual interest and Account B pays 5% annual interest. After one year, you have earned a total of $77.50 in interest. How
much did you invest in each account?

Answer by addingup(3677)

Let the amount you invest at 2% be x
Then, the amount invested at 5% will be 2000 - x
So:
0.02x + 0.05(2000 - x) = 77.50
0.02x + 100 - 0.05x = 77.50
-0.03x = -22.50
x = 750 this is the amount invested at 2%
2000 - 750 = the amount at 5%